# Unfolding the Literature: A Review of Robotic Cloth Manipulation

### v1

@@ 2.1 Textiles, Fabrics, and Cloths
The physical and mechanical properties of the final textile object are intrinsically tied to its manufacturing process. Yarns and threads, the building blocks of these textiles, are spun from various origins—animal, plant, mineral, synthetic, or blends thereof. Woven fabrics, produced by interlacing two sets of threads, are known for their hardness and non-elasticity, making them ideal for garments like shirts and jeans that retain creases. In contrast, knitted fabrics, created by interloping a single set of yarn, offer softness and elasticity in all directions and thus are used in wrinkle-resistant clothing such as t-shirts, which do not stretch uniformly in all directions to better conform to the body’s shape. Additionally, further production steps involve cutting and coloring fabrics to create clothing with various shapes and colored textures. A comprehensive exploration of fabric properties is given in [16].
@@ 2.2 Variations of Object Properties
One of the major challenges in robotic manipulation is designing planning and control algorithms that generalize or adapt to variations that will be encountered in real-world deployments [17]. In this review, we refer to generalization as the ability to apply the knowledge acquired from a specific set of textile properties to unseen variations of these properties. Adaptation is instead the ability to dynamically adjust strategies or models in response to changing conditions or novel variations in the object properties. Given the challenges of generalization and adaptation, it is crucial to first understand the variations in textile properties, which can be broadly classified into physical and mechanical properties. Fig. 2 shows an example of different variations of properties of textiles for four different objects.
@@ 2.2.1 Physical properties
The object’s physical properties are inherent characteristics and can be observed and measured without subjecting the material to external forces or manipulations. These properties describe how a material appears under static or non-changing conditions. Among these properties, Size and Shape are crucial as they influence the robot’s workspace, correlating with the object’s function and category. The Weight influences how the textile deforms under gravity. For heavy and large textiles such as blankets, robots may need bimanual systems to manipulate them. The Color affects perception-based algorithms, essential for identifying and tracking the textile during manipulation tasks [1]. Fabric Material composition, ranging from natural to synthetic fibers, dictates interaction forces between the textile and the robot, altering manipulation strategies [18]. The Construction Technique, referring to the knitting or weaving processes, directly impacts a fabric’s mechanical behavior by defining, for example, how thigh or loose the construction pattern is, thus influencing the object elasticity and rigidity [19].
@@ 3.3.1 Unmodeled Phenomenas
Robot simulators generally favour fast and visually correct behaviours at the expense of physical realism. Geometrically, textiles are often modelled as 2D objects without thickness. This requires extra care when the task concerns stacking or folding textiles of multiple layers. The mechanical modelling, be it physics-based or data-driven, rarely goes down to the basic building blocks of fabrics as laid out in Section 2.1. The simulation hence neglects aspects like fabric weave, thread count, as well as the response to real-world factors such as humidity, temperature, or wear and tear. In computer graphics, yarn-level simulation [49, 50] has been studied to account for warp and weft interactions. However, this has not yet been integrated into robotics modeling pipelines. The simplification of simulating effects in smaller granularity also impacts the fidelity of how a cloth interacts with external objects. Textiles likely exhibit more complex sliding behaviours along different directions due to how they are constructed. This breaks the popular isotropic friction cone model in many simulators. Missing yarn-level simulation also influences modeling the tear-up of a cloth. Current robot simulators often assume invariant topology and uniformity about how a cloth is constructed. Graphics research has demonstrated the possibility of capturing these nuances with advanced simulation [51, 49].
@@ 4 PROPERTIES PERCEPTION
Figure 3: Overview of sensors, properties, and actions in passive and interactive perception for textile manipulation. On the left, the figure presents a list of sensors, properties, and actions. The different colors of the sensors reflect to what extent the sensor has been explored for textile perception. On the right, it establishes connections between these categories based on the current literature, demonstrating how passive perception methods, using sensors like cameras and spectrometers, estimate properties that do not require physical interaction. In contrast, interactive perception integrates actions to improve the robot’s understanding of mechanical properties, leveraging specific actions coupled with sensor feedback to accurately perceive properties such as weight, stiffness, and elasticity.
@@ 4.1 Physical Properties Perception
As noted in Section 2.2.1, physical properties like size, shape, weight, color, fabric material, and construction technique can often be observed and measured without external manipulation. Particularly, size and shape can be effectively inferred by observing textiles spread over a flat surface and utilizing contour or segmentation techniques [59]. Data-driven techniques such as land-mark detection or end-to-end classifiers have also been explored for shape classification [60, 61]. However, textiles frequently arrive in deformed or crumpled states, complicating direct visual assessment. While some approaches have explored feature extraction processes to classify clothes from highly crumpled configurations [26], interactive perception is often necessary to enhance perceptual accuracy [62]. Lifting interactions, for example, are commonly employed to infer the shape of the textile better and facilitate its classification by garment type  [63, 64].
@@ 5.1.2 Data-driven Models
Data-driven models learn from data of interactions with the textile. There can be many state representations for data-driven models as laid out in Section 3.2. The choice of the state representation and model architecture can greatly affect the generalization ability of the model towards different properties such as the shape and texture of the textile. A particle-based state representation combined with graph neural networks [84, 85] has been shown to generalize better towards different shapes, geometries, and mechanical properties of the cloth compared to images [30] or latent-based state representations [83, 4], as the particle-based state representations align better with the underlying cloth physics. Using depth images [9] or point clouds [84, 103, 85] as the state representation also naturally makes the model invariant to the color and visual texture of the textiles. Due to the large amount of data needed for learning the model, most works use a simulator to generate the interaction data for learning the model [30, 83, 9, 84, 48, 85], with a few that does so in the real world [4].
@@ 5.2.1 Reinforcement Learning
To apply reinforcement learning (RL) to a textile manipulation problem, the manipulation problem needs to be formulated as a Markov Decision Process (MDP) or a Partially Observable MDP (POMDP). The key in the formulation is the design of the state, action, and reward space of the MDP. As in the case for model-based methods, common choices for state representations of textiles include manually defined features such as key points on the cloth [91, 96], or results from a perception system such as an image [47, 7, 88] or point cloud [20, 31] of the textile. Common actions for manipulating textiles include action primitives such as pick-and-placing [47], dragging [89], flinging [67], or raw actions that control the delta movement or velocity of the robot end-effector [88, 7, 31]. The reward function defines the desired outcome of the manipulation task based on the state of the environment.
Given the formulated MDP, reinforcement learning algorithms can be used for finding a policy that maximizes the expected accumulated rewards, thus learning the necessary manipulation skills [47, 7, 31, 88, 95]. Most works formulate the textile manipulation problem as a multi-step MDP [47, 31, 7, 88, 95] and learn a policy to maximize the accumulated reward, while some works [67, 89] formulate the problem into a bandit problem (an MDP with only 1 step), learn the reward function (usually in the form of a spatial-action value map [67, 28, 89]), and choose the best action as the one that maximizes the learned 1-step reward. Due to the large amount of data required by an RL algorithm, the need for a reward function, and the need to periodically reset the environment, such methods usually train a policy in a simulator and perform sim2real transfer [47, 7, 31, 91, 88, 95] with optional real-world fine-tuning [67]. A few works [86, 89, 28] directly train in the real world, where the reward can be automatically computed from real-world perception systems [89], and the environment can be automatically [28, 89], or manually [86] reset.
@@ 5.2.2 Imitation Learning
Task-wise, IL methods have been applied to cloth smoothing [94, 3, 92], folding [8, 93, 90, 87, 92], twisting [105], bed-making [99], and dressing [98]. Again, most works have only learned a single policy for one task [94, 8, 105, 90], with only a few that either learns a single backbone and different output branches on the same backbone [93], or a shared correspondence [92] for different tasks.
@@ 5.3 Heuristic-based Manipulation
The last discussed approach in textile manipulation is based on human-designed heuristic rules instead of learning the manipulation strategies from data. The heuristic rules vary depending on the target manipulation task. For example, for cloth smoothing, one commonly used rule is to detect the wrinkles of the cloth, and the manipulation action pulls the cloth in the perpendicular direction to the detected wrinkle direction [99]. For grasping, one rule is to use the position and orientation of a wrinkle to compute the target location and orientation of the gripper [29]. Another heuristic for a cloth article hanging in the air is to grasp key points that are identified based on the border geometry, e.g., corners for a towel, or sleeves for a t-shirt [101]. For folding, a heuristic folding motion can be achieved by leveraging gravity and moving the cloth such that the moved part is always vertical, which is called a “g-fold” [100]. For assistive dressing, common heuristic solution is to move the grasped garment along the forward direction of the human limb [53].
The generalization ability of such heuristically defined rules vary based on the target manipulation task and the assumptions made. Usually, these methods generalize well under the settings where the assumptions are satisfied, and tend to have limited generalization when the assumptions are broken. Due to the diverse properties and configurations a textile can have, it can be hard to design a heuristic rule that would generalize to every situation one might encounter when manipulating textiles.
@@ 6 BENCHMARKS AND DATASETS
Benchmarking robotic manipulation plays a crucial role in understanding methods’ capabilities and limitations enabling standardized comparison. Particularly benchmarking of cloth manipulation is limited due to the absence of objective and consistent evaluation processes and the limited research on how different textile materials, features and structures influence the performance of a method.
@@ 6.1 Benchmarks
Most benchmarks tackling cloth objects evaluate manipulation in simulated environments. Simulation benchmarks have the advantage that they can automatically generate many different conditions, enabling the evaluation of generalization. For instance, SoftGym [38] is a benchmark for deep reinforcement learning of deformable object manipulation that provides a set of simulated standardized environments and includes a set of cloth manipulation tasks. Assistive Gym [106] is a physics simulation framework for assistive tasks that includes an assistive dressing task with a hospital gown. However, due to the sim2real gap, it is necessary to have benchmarks for evaluating real-world applications with physical objects. Garcia-Camacho et al. [107] proposes benchmarks with real executions for three cloth manipulation tasks including protocols, qualitative evaluation metrics, and several complexity levels for each task based on the initial state of the cloth. Alternatively, Clark et al. [108] proposes four benchmark tasks for evaluating the performance of end-effectors in grasping clothing items, along with protocols to normalize starting crumpled configurations and metrics. Both works agree that an important aspect of benchmarks is the standardization of the objects. In addition, an effective way to measure the generalization of a method is through the use of a wide number of objects with varied properties. As it was done for rigid objects with the now widespread YCB object set [109], some efforts were put into replicating it for clothes [110], providing an extension of the YCB object set focusing on textile objects. It includes a wide variety of cloth household objects with benchmarking guidelines for its use. This object set was distributed among research institutes and the participants of the cloth manipulation and perception competition organized in IROS’22 and ICRA’23 to extend its usage, a requirement for a benchmark to be useful. However, there is an important issue in defining standardized textile object sets given by stock continuity since the fast-fashion industry changes stock each season, preventing the maintenance of the same objects for extended time periods. To solve this issue, a method for building comparable textile object sets across different publications has been proposed in [23]. It proposes a framework to characterize textile objects through the measurement of their physical and mechanical properties, enabling the quantification of variability on the textile properties listed in Section 2 of a given set of objects. With this characterization, the generalization of a method can be quantified based on the amount of variation that the objects offer. This idea goes in line with the one proposed for rigid objects in [111], where generalization is measured similarly but also adds other aspects of variability in background, table color, etc. Variability in the starting configurations and scene properties of color and light should also be considered for cloth manipulation.
@@ 6.2 Datasets
Unlike benchmarks, datasets serve the purpose of providing data for designing or learning a task. The extent of generalization depends on the variation covered by the dataset. Datasets are often task-specific, ensuring that the data is relevant both for training a system and evaluating its performance. In cloth manipulation, the most common datasets include simulated 3D models [112, 113], RGB images [61, 114, 115] or depth [89].
The largest existing datasets are for classic vision problems like cloth classification and landmark point detection, with datasets such as ClothesNet [112] in simulation or DeepFashion [61] with real images and segmentation of people wearing all type of clothes with datasets such as CLOTH3D [113] in simulation or [116] with real images. Surface reconstruction is another classic perception problem that has been applied to cloth [114]. For cloth classification or landmark detection, images contain annotations of cloth type and location of landmarks, but usually, images come from the fashion industry, and so clothes are either flatted, hung on hangers, or worn by humans. For surface reconstruction, annotations need to have realistic images with the real mesh of the object, and therefore, existing datasets are all in simulation and subject to the sim2real gap discussed in Section 3.3.
@@
One of the challenges for datasets in cloth manipulation is labelling the ground truth of the deformation state in real images. Only a few datasets exist with point clouds with labels of what points correspond to corners or edges during a manipulation [120]. So far, to encode interactions, datasets use RGB-D images where the action is annotated as a pick-up pixel point and a direction of motion in the image [89], with the limitation of only representing close-to-planar configurations. More complex actions appeared lately with Visual Language Models where a sequence of images is linked to a sequence of positions of the end-effector [121], with corresponding datasets.
@@ 7 APPLICATIONS AREAS
We outline the applications and challenges of manipulating deformable textile objects in scenarios requiring robots to adapt and generalize to their varying properties. Understanding the physical properties of textiles is crucial for a wide range of tasks in diverse sectors such as household chores, healthcare, and the fashion industry. Table 2 provides a (non-exhaustive) overview of work categories in these sectors, organized by the frequency at which they are addressed in the literature. Tasks such as folding, smoothing, and dressing receive frequent attention from the community, whereas tasks like buttoning, dyeing, and washing remain rather underexplored. In what follows, we will discuss in detail these tasks and the requirements concerning variations of textile properties.
@@
Frequency Household Healthcare Fashion Industry

### v2

@@ 2.1 Textiles, Fabrics, and Cloths
The physical and mechanical properties of the final textile object are intrinsically tied to its manufacturing process. Yarns and threads, the building blocks of these textiles, are spun from various origins—animal, plant, mineral, synthetic, or blends thereof. Woven fabrics, produced by interlacing two sets of threads, are known for their hardness and non-elasticity, making them ideal for garments like shirts and jeans that retain creases. In contrast, knitted fabrics, created by interloping a single set of yarn, offer softness and elasticity in all directions and thus are used in wrinkle-resistant clothing such as t-shirts, which stretch uniformly in all directions to better conform to the body’s shape. Additionally, further production steps involve cutting and coloring fabrics to create clothing with various shapes and colored textures. A comprehensive exploration of fabric properties is given in [16].
@@ 2.2 Variations of Object Properties
One of the major challenges in robotic manipulation is designing planning and control algorithms that generalize or adapt to variations that will be encountered in real-world deployments [17]. In this review, we refer to generalization as the ability to apply the knowledge acquired from a specific set of textile properties to unseen variations of these properties. Adaptation is instead the ability to dynamically adjust strategies or models in response to changing conditions or novel variations in the object properties. Given the challenges of generalization and adaptation, it is crucial to first understand the variations in textile properties, which can be broadly classified into physical and mechanical properties. Fig. 2 shows an example of different variations of textile properties for four different objects.
@@ 2.2.1 Physical properties
The object’s physical properties are inherent characteristics and can be observed and measured without subjecting the material to external forces or manipulations. These properties describe how a material appears under static or non-changing conditions. Among these properties, Size and Shape are crucial as they influence the robot’s workspace, correlating with the object’s function and category. The Weight influences how the textile deforms under gravity. For heavy and large textiles such as blankets, robots may need bimanual systems to manipulate them. The Color affects perception-based algorithms, essential for identifying and tracking the textile during manipulation tasks [1]. Fabric Material composition, ranging from natural to synthetic fibers, dictates interaction forces between the textile and the robot, altering manipulation strategies [18]. The Construction Technique, referring to the knitting or weaving processes, directly impacts a fabric’s mechanical behavior by defining, for example, how tight or loose the construction pattern is, thus influencing the object elasticity and rigidity [19].
@@ 3.3.1 Unmodeled Phenomenas
Robot simulators generally favour fast and visually correct behaviours at the expense of physical realism. Geometrically, textiles are often modelled as 2D objects without thickness. This requires extra care when the task concerns stacking or folding textiles of multiple layers. The mechanical modelling, be it physics-based or data-driven, rarely goes down to the basic building blocks of fabrics as laid out in Section 2.1. The simulation hence neglects aspects like fabric weave, thread count, as well as the response to real-world factors such as humidity, temperature, or wear and tear. In computer graphics, yarn-level simulation [49, 50] has been studied to account for warp and weft interactions. However, this has not yet been integrated into robotics modeling pipelines. The simplification of simulating effects in smaller granularity also impacts the fidelity of how a cloth interacts with external objects. Textiles likely exhibit more complex sliding behaviours along different directions due to how they are constructed. This breaks the popular isotropic friction cone model in many simulators. Missing yarn-level simulation also influences modeling the tear-up of a cloth. Current robot simulators often assume invariant topology and uniformity about how a cloth is constructed. Graphics research has demonstrated the possibility of capturing these nuances with advanced simulation [49, 51]. Furthermore, another phenomena that is rarely included in cloth models is the aerodynamics of the textiles. Since most garments are very light, even in the absence of wind, the air that surrounds them has a critical impact on highly dynamic cloth motions [52]. Much research is still needed in order to have realistic, yet efficient physical models of cloth aerodynamics.
@@ 4 PROPERTIES PERCEPTION
Figure 3: Overview of sensors, properties, and actions in passive and interactive perception for textile manipulation. On the left, the figure presents a list of sensors, properties, and actions. LR and HR for tactile sensors stand for low resolution and high resolution, respectively. The variation in opacity of the boxes reflects to what extent the sensor or the property has been explored for textile perception. On the right, the figure shows connections between sensors, properties, and actions based on the current literature. This side of the figure showcases how passive perception methods, using sensors like cameras and spectrometers, estimate properties that do not require physical interaction. In contrast, interactive perception integrates actions to improve the robot’s understanding of mechanical properties, leveraging specific actions coupled with sensor feedback to accurately perceive properties such as weight, stiffness, and elasticity.
@@ 4.1 Physical Properties Perception
As noted in Section 2.2.1, physical properties like size, shape, weight, color, fabric material, and construction technique can often be observed and measured without external manipulation. Particularly, size and shape can be effectively inferred by observing textiles spread over a flat surface and utilizing contour or segmentation techniques [60]. Data-driven techniques such as landmark detection or end-to-end classifiers have also been explored for shape classification [61, 62]. However, textiles frequently arrive in deformed or crumpled states, complicating direct visual assessment. While some approaches have explored feature extraction processes to classify clothes from highly crumpled configurations [26], interactive perception is often necessary to enhance perceptual accuracy [63]. Lifting interactions, for example, are commonly employed to infer the shape of the textile better and facilitate its classification by garment type  [64, 65].
@@ 5.1.2 Data-driven Models
Data-driven models learn from data about interactions with textiles. There can be many state representations for data-driven models as laid out in Section 3.2. The choice of the state representation and model architecture can greatly affect the generalization ability of the model towards different properties, such as the shape and texture of the textile. A particle-based state representation combined with graph neural networks [85, 86] has been shown to generalize better towards different shapes, geometries, and mechanical properties of the cloth compared to images [30] or latent-based state representations [4, 84], as the particle-based state representations align better with the underlying cloth physics. Using depth images [9] or point clouds [85, 86, 104] as the state representation also naturally makes the model invariant to the color and visual texture of the textiles. Due to the large amount of data needed for learning the model, most works use a simulator to generate the interaction data for learning the model [9, 30, 48, 84, 85, 86], with a few that does so in the real world [4].
@@ 5.2.1 Reinforcement Learning
To apply reinforcement learning (RL) to a textile manipulation problem, the manipulation problem needs to be formulated as a Markov Decision Process (MDP) or a Partially Observable MDP (POMDP). The key in the formulation is the design of the state, action, and reward space of the MDP. As in the case of model-based methods, common choices for state representations of textiles include manually defined features such as key points on the cloth [92, 97], or results from a perception system such as an image [7, 47, 89] or point cloud [20, 31] of the textile. Common actions for manipulating textiles include action primitives such as pick-and-placing [47], dragging [90], flinging [68], or raw actions that control the delta movement or velocity of the robot end-effector [7, 31, 89]. The reward function defines the desired outcome of the manipulation task based on the state of the environment.
Given the formulated MDP, reinforcement learning algorithms can be used to find a policy that maximizes the expected accumulated rewards, thus learning the necessary manipulation skills. Most works formulate the textile manipulation problem as a multi-step MDP [7, 31, 47, 89, 96] and learn a policy to maximize the accumulated reward, while some works [68, 90] formulate the problem into a bandit problem (an MDP with only 1 step), learn the reward function (usually in the form of a spatial-action value map [28, 68, 90]), and choose the best action as the one that maximizes the learned 1-step reward. Due to the large amount of data required by an RL algorithm, the need for a reward function, and the need to periodically reset the environment, such methods usually train a policy in a simulator and perform sim2real transfer [7, 31, 47, 89, 92, 96] with optional real-world fine-tuning [68]. A few works [28, 87, 90] directly train in the real world, where the reward can be automatically computed from real-world perception systems [90], and the environment can be automatically [28, 90], or manually [87] reset.
@@ 5.2.2 Imitation Learning
Task-wise, IL methods have been applied to cloth smoothing [3, 93, 95], folding [8, 88, 91, 93, 94], twisting [106], bed-making [100], and dressing [99]. Again, most works have only learned a single policy for one task [8, 91, 95, 106], with only a few that either learn a single backbone and different output branches on the same backbone [94], or a shared correspondence [93] for different tasks.
@@ 5.3 Heuristic-based Manipulation
The last discussed approach in textile manipulation is based on human-designed heuristic rules instead of learning the manipulation strategies from data. The heuristic rules vary depending on the target manipulation task. For example, for cloth smoothing, one commonly used rule is to detect the wrinkles of the cloth, and the manipulation action pulls the cloth in the perpendicular direction to the detected wrinkle direction [100]. For grasping, one rule is to use the position and orientation of a wrinkle to compute the target location and orientation of the gripper [29]. Another heuristic for a cloth article hanging in the air is to grasp key points that are identified based on the border geometry, e.g., corners for a towel or sleeves for a t-shirt [102], and use force sensors to trace the edge and get to the opposite corner [107]. For folding, a heuristic folding motion can be achieved by leveraging gravity and moving the cloth such that the moved part is always vertical, which is called a “g-fold” [101]. For assistive dressing, a common heuristic solution is to move the grasped garment along the forward direction of the human limb [54].
The generalization ability of such heuristically defined rules varies based on the target manipulation task and the assumptions made. Usually, these methods generalize well under the settings where the assumptions are satisfied and tend to have limited generalization when the assumptions are broken. Due to the diverse properties and configurations a textile can have, it can be hard to design a heuristic rule that would generalize to every situation one might encounter when manipulating textiles.
@@ 6 BENCHMARKS AND DATASETS
Benchmarking robotic manipulation plays a crucial role in understanding methods’ capabilities and limitations, enabling standardized comparison. Particularly benchmarking of cloth manipulation is limited due to the absence of objective and consistent evaluation processes and the limited research on how different textiles influence the performance of a method.
@@ 6.1 Benchmarks
Most benchmarks tackling cloth objects evaluate manipulation in simulated environments, since they have the advantage of automatically generating various conditions to evaluate generalization. For instance, SoftGym [38] for deformable object manipulation and Assistive Gym [108] for assistive tasks, are simulation benchmarks for reinforcement learning that provides a set of simulated standardized environments. However, due to the sim2real gap, it is necessary to have benchmarks for evaluating real-world applications with physical objects. Garcia-Camacho et al. [109] proposes benchmarks with real executions for three cloth manipulation tasks including protocols, qualitative evaluation metrics, and several complexity levels based on the initial state of the cloth. Alternatively, Clark et al. [110] proposes four benchmark tasks for evaluating the performance of end-effectors in grasping clothing items, along with protocols to normalize crumpled configurations and metrics.
An effective way to measure the generalization of a method is through the use of a wide number of objects with varied properties. As it was done for rigid objects with the now widespread YCB object set [111], an extension focusing on textile objects is provided in [112], which was distributed among the participants of the cloth manipulation and perception competition [113]. It includes a wide variety of cloth household objects with benchmarking guidelines for its use. However, an important issue in defining standardized textile object sets is stock continuity, preventing the maintenance of the same objects for extended time periods. To solve this issue, a method for building comparable textile object sets across different publications has been proposed in [23]. It proposes a framework to characterize textile objects, enabling the quantification of variability on the textile properties listed in Section 2 of a given set of objects. With this characterization, the generalization of a method can be quantified based on the amount of variation that the objects offer. This idea goes in line with the one proposed for rigid objects in [114], where generalization is measured similarly but also adds other aspects of variability in background, table color, etc.
@@ 6.2 Datasets
Unlike benchmarks, datasets serve the purpose of providing data for designing or learning a task. The extent of generalization depends on the variation covered by the dataset. Datasets are often task-specific, ensuring that the data is relevant both for training a system and evaluating its performance. In cloth manipulation, the most common datasets include simulated 3D models [115, 116], RGB images [62, 117, 118] or depth [90, 119].
The largest existing datasets are for classic vision problems like cloth classification and landmark point detection, with datasets such as ClothesNet [115] in simulation or DeepFashion [62] with real images. Datasets for segmentation of people wearing clothes include CLOTH3D [116] in simulation or [120] with real images. Surface reconstruction is another classic perception problem applied to cloth [117]. For cloth classification or landmark detection, images contain annotations of cloth type and location of landmarks, but usually, images come from the fashion industry, and so clothes are either flatted, hung on hangers, or worn by humans. For surface reconstruction, annotations need to have realistic images with the real mesh of the object, and therefore, existing datasets are all in simulation and subject to the sim2real gap discussed in Section 3.3.
@@
One of the challenges for datasets in cloth manipulation is labelling the ground truth of the deformation state in real images. Only a few datasets exist with depth or point clouds [119] and even less with labels of what points correspond to corners or edges during a manipulation [124]. So far, to encode interactions, some datasets use RGB-D images where the action is annotated as a pick-up pixel point and a direction of motion in the image [90], with the limitation of only representing close-to-planar configurations. More complex actions appeared lately with Visual Language Models where a sequence of images is linked to a sequence of positions of the end-effector [125], with corresponding datasets.
@@ 7 APPLICATIONS AREAS
We outline the applications and challenges of manipulating deformable textile objects in scenarios requiring robots to adapt and generalize to their varying properties. Understanding the physical properties of textiles is crucial for a wide range of tasks in diverse sectors such as household chores, healthcare, and the textile industry. Table 2 provides a (non-exhaustive) overview of work categories in these sectors, organized by the frequency at which they are addressed in the literature. Tasks such as folding, smoothing, and dressing receive frequent attention from the community, whereas tasks like buttoning, dyeing, and washing remain rather underexplored. In what follows, we will discuss in detail these tasks and the requirements concerning variations of textile properties.
@@
Frequency Household Healthcare Textile Industry